provide a detailed cost benefit analysis of taking online accelerated microeconomics class during your winter holidays. Try to quantify some of the economic terms that you will use in your own analysis, to the extent possible.

Book cost : $145

3 unit credit class tution fee $160

You can create an example of the costs and benefit the above situation hypothetically.
Answer #1

The costs involved here are cost of the book and the tuition fee. So total cost = 145+160 = $305

There will also be an opportunity cost of losing out on leisure time of winter holidays. It should be noted that opportunity costs are those costs that are involved when a particular course of action is taken. The opportunity cost will also involve the pay that you would have earned if you had taken a part time job during the winter holiday. Suppose that the opportunity costs are $300 in total. Thus your total costs = $145+160+300 = $605.

In terms of benefits you will enhance and augment your skills through the online accelerated microeconomics class and this will increase your demand in the job market in the long run as your skill sets will be at a higher level and hence you will be more relevant as a candidate for a job related to economics. The job market is a competitive market and by taking the online course you will be able to ensure that you are more sellable as a candidate compared to other candidates. It will also help in earning you a better salary.

Suppose that the present value of all benefits (as mentioned above) comes to a total of $1,000. Hence your benefit cost ratio will be = $1000/605 = 1.65

As the ratio>1 you can see that it will be in your interest to take the online class.